OpenStudy (anonymous):

If 4 is subtracted from twice a number, the result is 12 less than the number. A. 2x – 4 = x + 12 x = 16 B. 2x + 4 = x – 12 x = –16 C. 2x + 4 = x + 12 x = 8 D. 2x – 4 = x – 12 x = –8

mathslover (mathslover):

Let the number be x. \(\color{Red}{\textbf{Twice the number = 2x.}} \) "4 is subtracted from \(\color{red}{\bf{twice~ a~ number}}\)" => "4 is subtracted from \(\color{red}{\bf{2x}}\) That means, 2x - 4 As the result is 12 less than number, that is : x - 12. Therefore , 2x - 4 = x-12

OpenStudy (anonymous):

so x equals -16 or -8?

mathslover (mathslover):

-8. \(2x -4 = x-12\) \(2x - x = -12 + 4\) \(x = -8\)